I hope you can help me with this problem:
A ski trip at the school has been arranged. There are 30 students that have paid for the trip and 6 parnts that have volunteered to chaperone. To transport the students and parents easily they are to be divided into two groups. One group has 10 students and 2 parents, and the other group has 20 students and 4 parents.
A) How many different groups of ten students can be formed?
B) How many diffeent groups consisting of 10 students and 2 parents are posible?
C) Kelly is one of the students going on the trip, and Kelly's mother is a chaperone. Kelly's mother would prefer to be in the smaller group and not in the same group as Kelly. If this wish is honoured, how many ways can the smaller group and its chaperone be chosen?

Answer by stanbon(75887)

A ski trip at the school has been arranged. There are 30 students that have paid for the trip and 6 parnts that have volunteered to chaperone. To transport the students and parents easily they are to be divided into two groups. One group has 10 students and 2 parents, and the other group has 20 students and 4 parents.
A) How many different groups of ten students can be formed?
30C10= 30,045,015
----------------------
B) How many diffeent groups consisting of 10 students and 2 parents are posible?
(30C10)*(6C2)=450,675,225
------------------------
C) Kelly is one of the students going on the trip, and Kelly's mother is a chaperone. Kelly's mother would prefer to be in the smaller group and not in the same group as Kelly. If this wish is honoured, how many ways can the smaller group and its chaperone be chosen?
# of ways to choose 2 parents if Kelly's mom is one of them= 1*5=5
# of ways to choose 10 students if Kelly is excluded= 29C10
# ways to meet both conditions: 5(29C10)=5*20,030,010=100,150,050
Cheers,
Stan H.